OpenStudy (anonymous):

The pressure of a gas p(v) varies inversely with the volume of the gas v. The pressure of a gas measures 25 kg/cm2 when its volume is 200 cm3. Which equation can be used to find the pressure of the gas when the volume is changed?

OpenStudy (anonymous):

p(v)=8/v p(v)=8v p(v)=5000/v p(v)=5000v

OpenStudy (anonymous):

Please help me :(

OpenStudy (anonymous):

pv = nrt?

OpenStudy (anonymous):

if presure increase, then volume has to decrease

OpenStudy (anonymous):

I dont know how I'd find it..

OpenStudy (anonymous):

Would it be the first one?

OpenStudy (anonymous):

or the 3rd one?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

*sorry, i'm taking so long, im getting spread pretty thin between everyone ^_^

OpenStudy (anonymous):

its okay!

OpenStudy (anonymous):

if P(v) = 25, and v=200, which equation is true? p(v)=8/v p(v)=8v p(v)=5000/v p(v)=5000v check: p(v)=8/v 25 = 8/200 [FALSE] check the rest of the choices like i just checked the first one

OpenStudy (anonymous):

u get the answer?

OpenStudy (anonymous):

I think. D?...

OpenStudy (anonymous):

D or C? @DemolisionWolf

OpenStudy (anonymous):

p(v)=5000/v 25 = 5000/200 [TRUE] p(v)=5000v 25 = 5000(200) [FALSE]

OpenStudy (anonymous):

have a nice night ^_^

OpenStudy (anonymous):

You;re amazing! Thanks! you too! @DemolisionWolf